# Hadronic tensor molecule $B_{c}^{\ast+}B_{c}^{\ast-}$

### v1

@@ Abstract
Mass and full decay width of the hadronic tensor molecule $\mathcal{M}_{\mathrm{T}}=B_{c}^{\ast+}B_{c}^{\ast-}$ are examined in the framework of QCD sum rule method. To find the mass $m$ and current coupling $\Lambda$ of this state, we use the two-point sum rules. The result $(12.87\pm 0.08)~\mathrm{GeV}$ for $m$ indicates that $\mathcal{M}_{\mathrm{T}}$ can easily decay to pairs of $J/\psi\Upsilon$, $\eta_{b}\eta_{c}$, $B_{c}^{+}B_{c}^{-}$, and $B_{c}^{\ast+}B_{c}^{\ast-}$ mesons. Kinematically permitted ways for $\mathcal{M}_{\mathrm{T}}$ to transform to conventional mesons include also processes $\mathcal{M}_{\mathrm{T}}\to D^{(\ast)+}D^{(\ast)-}$, $D^{(\ast)0}\overline{D}^{(\ast)0}$, and $D_{s}^{(\ast)+}D_{s}^{(\ast)-}$ triggered by annihilation of $b\overline{b}$ quarks in the molecule. The decays to meson pairs $B^{(\ast)+}B^{(\ast)-}$, $B^{(\ast)0}\overline{B}^{(\ast)0}$, and $B_{s}^{(\ast)0}\overline{B}_{s}^{(\ast)0}$ generated by $c\overline{c}\to$ light quarks are also among possible channels of the hadronic molecule $\mathcal{M}_{\mathrm{T}}$. Technical tools of the three-point sum rule approach are applied to compute partial widths of these decays. Our results for the mass and width $\Gamma_{\mathcal{M}_{\mathrm{T}}}=(154\pm 19)~\mathrm{MeV}$ of the tensor molecule $\mathcal{M}_{\mathrm{T}}$ are important for experimental studies of fully heavy exotic structures.
@@ I Introduction
The diquark-antidiquark structures $bc\overline{b}\overline{c}$ with numerous quantum numbers were intensively investigated in the literature Faustov:2022mvs ; Wu:2016vtq ; Liu:2019zuc ; Chen:2019vrj ; Bedolla:2019zwg ; Cordillo:2020sgc ; Weng:2020jao ; Yang:2021zrc ; Hoffer:2024alv . In these articles the authors concentrated mainly on their masses and computed this parameter by employing different methods. It should be noted that the decays of the tetraquarks $bc\overline{b}\overline{c}$, as usual, remained beyond the scope of these studies. In our publications Agaev:2024wvp ; Agaev:2024mng ; Agaev:2024qbh we considered the diquark-antidiquark states $bc\overline{b}\overline{c}$ with spin-parities $J^{\mathrm{P}}=0^{+}$, $1^{+}$, $2^{+}$ and calculated their masses and widths.
@@ II Spectroscopic parameters $m$ and $\Lambda$ of the molecule $\mathcal{M}_{\mathrm{T}}$
has to satisfy $\mathrm{PC}\geq 0.5$ which leads to its dominance in $m$ and $\Lambda$. The correlation function contains only the dimension-$4$ term $\Pi^{\mathrm{Dim4}}(M^{2},s_{0})$, therefore we demand fulfilment of the constraint $|\Pi^{\mathrm{Dim4}}(M^{2},s_{0})|\leq 0.05|\Pi(M^{2},s_{0})|$: This ensures the convergence of $\mathrm{OPE}$. The maximum of the Borel parameter is fixed by Eq. (11), whereas convergence of $\mathrm{OPE}$ permits us to obtain its minimal value.
@@
The results in Eq. (13) amount to SR predictions at the point $M^{2}=9.5~\mathrm{GeV}^{2}$ and $s_{0}=185~\mathrm{GeV}^{2}$, where $\mathrm{PC}\approx 0.59$. This guaranties the prevalence of $\mathrm{PC}$ in the quantities $m$ and $\Lambda$. The main sources of ambiguities in Eq. (13) are parameters $M^{2}$ and $s_{0}$. Relevant errors constitutes $\pm 0.6\%$ of the mass $m$, which proves reliability of the extracted result. Such accuracy is achieved due to the sum rule for $m$, Eq. (7), which is a ratio of correlators. Consequently, variations of these functions due to changes of $M^{2}$, and $s_{0}$ recompense in $m$ and stabilize the result. Ambiguities of the current coupling $\Lambda$ are equal to $\pm 11.6\%$ remaining nevertheless inside borders allowable for SR studies. In Fig. 2, we depict $m$ as a function of $M^{2}$ and $s_{0}$.
@@ III.1 Decay $\mathcal{M}_{\mathrm{T}}\rightarrow J/\psi\Upsilon$
$\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})=\int_{4\mathcal{M}^{2}}^{s_{0}}ds\int_{4m_{b}^{2}}^{s_{0}^{\prime}}ds^{\prime}\rho_{1}(s,s^{\prime},q^{2})$
@@
$g_{1}\equiv\mathcal{G}_{1}(-m_{J/\psi}^{2})=(2.9\pm 0.5)\times 10^{-1}\ \mathrm{GeV}^{-1}.$ (26)
@@
The errors originate from the ambiguities in the strong coupling $g_{1}$ and those of the particles’ masses in Eq. (27)
Figure 3: QCD data and fit function for $g_{1}(Q^{2})$. The diamond fixes the point $Q^{2}=-m_{J/\psi}^{2}$ where $g_{1}$ has been estimated.
@@ V Decays generated by $b\overline{b}$ annihilations
These modes are kinematically possible channels of the molecule $\mathcal{M}_{\mathrm{T}}$ to conventional particles. We investigate these decays also in the three-point SR framework. Here, to calculate the correlation functions which contain vacuum matrix element $\langle\overline{b}b\rangle$ of $b$ quark, we replace it with the known gluon condensate $\langle\alpha_{s}G^{2}/\pi\rangle$ Agaev:2023ara .
@@ VI Decays into $B$ meson pairs
Decays of the hadronic molecule $\mathcal{M}_{\mathrm{T}}$ into pairs of the $B$ and $B_{s}$ mesons are investigated by the same manner. These decays are generated by annihilation in the molecule of $c\overline{c}$ quarks which lead to creations of $BB$ and $B_{s}B_{s}$ mesons with correct charges and quantum numbers. It is easy to see that $\mathcal{M}_{\mathrm{T}}\rightarrow B^{(\ast)+}B^{(\ast)-}$, $B^{(\ast)0}\overline{B}^{(\ast)0}$, and $B_{s}^{0}\overline{B}_{s}^{0}$ are allowed decay channels of the hadronic molecule $\mathcal{M}_{\mathrm{T}}$.
@@
Information on the partial widths of the channels considered in last three sections permit us to estimate the full width of the hadronic molecule

### v2

@@ Abstract
Mass and full decay width of the hadronic tensor molecule $\mathcal{M}_{\mathrm{T}}=B_{c}^{\ast+}B_{c}^{\ast-}$ with spin-parities $J^{\mathrm{PC}}=2^{++}$ are examined in the framework of QCD sum rule method. To find the mass $m$ and current coupling $\Lambda$ of this state, we use the two-point sum rules. The result $(12.87\pm 0.08)~\mathrm{\ GeV}$ for $m$ indicates that $\mathcal{M}_{\mathrm{T}}$ can easily decay to pairs of $J/\psi\Upsilon$, $\eta_{b}\eta_{c}$, $B_{c}^{+}B_{c}^{-}$, and $B_{c}^{\ast+}B_{c}^{\ast-}$ mesons. Kinematically permitted ways for $\mathcal{M}_{\mathrm{T}}$ to transform to conventional mesons include also processes $\mathcal{M}_{\mathrm{T}}\to D^{(\ast)+}D^{(\ast)-}$, $D^{(\ast)0}\overline{D}^{(\ast)0}$, and $D_{s}^{(\ast)+}D_{s}^{(\ast)-}$ triggered by annihilation of $b\overline{b}$ quarks in the molecule. The decays to meson pairs $B^{(\ast)+}B^{(\ast)-}$, $B^{(\ast)0}\overline{B}^{(\ast)0}$ , and $B_{s}^{(\ast)0}\overline{B}_{s}^{(\ast)0}$ generated by $c\overline{c}\to$ light quarks are also among possible channels of the hadronic molecule $\mathcal{M}_{\mathrm{T}}$. Technical tools of the three-point sum rule approach are applied to compute partial widths of these decays. Our results for the mass and width $\Gamma_{\mathcal{M}_{\mathrm{T}}}=(154\pm 19)~\mathrm{MeV}$ of the tensor molecule $\mathcal{M}_{\mathrm{T}}$ are important for experimental studies of fully heavy exotic structures.
@@ I Introduction
The diquark-antidiquark structures $bc\overline{b}\overline{c}$ with numerous quantum numbers were intensively investigated in the literature Faustov:2022mvs ; Wu:2016vtq ; Liu:2019zuc ; Chen:2019vrj ; Bedolla:2019zwg ; Cordillo:2020sgc ; Weng:2020jao ; Yang:2021zrc ; Hoffer:2024alv . In these articles the authors concentrated mainly on their masses and computed them by employing different methods. It should be noted that the decays of the tetraquarks $bc\overline{b}\overline{c}$, as usual, remained beyond the scope of these studies. In our publications Agaev:2024wvp ; Agaev:2024mng ; Agaev:2024qbh we considered the diquark-antidiquark states $bc\overline{b}\overline{c}$ with spin-parities $J^{\mathrm{P}}=0^{+}$, $1^{+}$, $2^{+}$ and calculated their masses and widths.
@@ II Spectroscopic parameters $m$ and $\Lambda$ of the molecule $\mathcal{M}_{\mathrm{T}}$
has to satisfy $\mathrm{PC}>0.5$ which leads to its dominance in $m$ and $\Lambda$. The correlation function contains only the dimension-$4$ term $\Pi^{\mathrm{Dim4}}(M^{2},s_{0})$, therefore we demand fulfilment of the constraint $|\Pi^{\mathrm{Dim4}}(M^{2},s_{0})|\leq 0.05|\Pi(M^{2},s_{0})|$: This ensures the convergence of $\mathrm{OPE}$. The maximum of the Borel parameter is fixed by Eq. (11), whereas convergence of $\mathrm{OPE}$ permits us to obtain its minimal value.
@@
The results in Eq. (13) amount to SR predictions at the point $M^{2}=9.5~\mathrm{GeV}^{2}$ and $s_{0}=185~\mathrm{GeV}^{2}$, where $\mathrm{PC}\approx 0.59$. This guaranties prevalence of $\mathrm{PC}$ in the quantities $m$ and $\Lambda$. The main sources of ambiguities in Eq. (13) are parameters $M^{2}$ and $s_{0}$: Errors generated by uncertainties in the quark masses and gluon condensate are negligibly small. Relevant errors constitutes $\pm 0.6\%$ of the mass $m$, which proves reliability of the extracted result. Such accuracy is achieved due to the sum rule for $m$, Eq. (7), which is a ratio of correlators. Consequently, variations of these functions due to changes of $M^{2}$, and $s_{0}$ recompense in $m$ and stabilize the result. Ambiguities of the current coupling $\Lambda$ are equal to $\pm 11.6\%$ remaining nevertheless inside borders allowable for SR studies. In Fig. 2, we depict $m$ as a function of $M^{2}$ and $s_{0}$.
@@ III.1 Decay $\mathcal{M}_{\mathrm{T}}\rightarrow J/\psi\Upsilon$
$\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})=\int_{4(m_{b}+m_{c})^{2}}^{s_{0}}ds\int_{4m_{b}^{2}}^{s_{0}^{\prime}}ds^{\prime}\rho_{1}(s,s^{\prime},q^{2})$
@@
$g_{1}\equiv\mathcal{G}_{1}(-m_{J/\psi}^{2})=(2.86\pm 0.50)\times 10^{-1}\ \mathrm{GeV}^{-1}.$ (26)
@@
The errors above originate from the ambiguities in the strong coupling $g_{1}$ and those of the particles’ masses in Eq. (27). A potential source of theoretical errors is also a choice of the interpolation function $\mathcal{G}_{i}(Q^{2})$ in the form of Eq. (24). It should be emphasized that alternative forms for this function, in general, may lead to deviations from Eq. (26), but as usual they are very small and can be neglected. In fact, let us consider the fit function in the following analytical form
$\overline{\mathcal{G}}_{1}(Q^{2})=\frac{\overline{\mathcal{G}}_{i}^{0}}{\left(1-\frac{Q^{2}}{m^{2}}\right)\left(1-\sigma_{1}\frac{Q^{2}}{m^{2}}+\sigma_{2}\left(\frac{Q^{2}}{m^{2}}\right)^{2}\right)},$ (31)
where $\overline{\mathcal{G}}_{1}^{0}$, $\sigma_{1}$ and $\sigma_{2}$ are fitting constants.
Then, using the QCD data and Eq. (31), one finds the parameters $\overline{\mathcal{G}}_{1}^{0}=0.319$, $\sigma_{1}=1.244$ and $\sigma_{2}=0.402$ in $\overline{\mathcal{G}}_{1}(Q^{2})$. The function $\overline{\mathcal{G}}_{1}(Q^{2})$ is also depicted in Fig. 3, where one can see its nice agreement with SR data. Having used this function, we get $\overline{g}_{1}=0.281~\mathrm{GeV}^{-1}$. As is seen, a difference $g_{1}-\overline{g}_{1}=0.005$ is an order of magnitude smaller than ambiguities $\pm 0.05$ of $g_{1}$ from Eq. (26).Therefore, throughout this work we employ Eq. (24) and neglect small effects due to alternative extrapolating functions.
Figure 3: QCD data and fit functions for $g_{1}(Q^{2})$ (solid line) and $\overline{g}_{1}(Q^{2})$ (dashed line). The diamond fixes the point $Q^{2}=-m_{J/\psi}^{2}$ where $g_{1}$ and $\overline{g}_{1}$ have been estimated.
@@ V Decays generated by $b\overline{b}$ annihilations
These modes are kinematically possible channels of the molecule $\mathcal{M}_{\mathrm{T}}$ to conventional particles. We investigate these decays also in the three-point SR method. Here, to calculate the correlation functions which contain vacuum matrix element $\langle\overline{b}b\rangle$ of $b$ quark, we replace it with the known gluon condensate $\langle\alpha_{s}G^{2}/\pi\rangle$ Agaev:2023ara . But implementation of annihilation processes into the three-point SR framework does not imply usage of additional free parameters.
@@ VI Decays into $B$ meson pairs
Decays of the hadronic molecule $\mathcal{M}_{\mathrm{T}}$ into pairs of the $B$ and $B_{s}$ mesons are investigated by the same manner. These decays are generated by $c\overline{c}$ annihilation which lead to creations of $BB$ and $B_{s}B_{s}$ mesons with correct charges and quantum numbers. It is easy to see that $\mathcal{M}_{\mathrm{T}}\rightarrow B^{(\ast)+}B^{(\ast)-}$, $B^{(\ast)0}\overline{B}^{(\ast)0}$, and $B_{s}^{0}\overline{B}_{s}^{0}$ are allowed decay channels of the hadronic molecule $\mathcal{M}_{\mathrm{T}}$.
@@
Information on the partial widths of the channels considered in last three sections permits us to estimate the full width of the hadronic molecule